# Exact Hull Reformulation for Quadratically Constrained Generalized Disjunctive Programs

### v1

@@ 6 Computational experiments
The primary objective of the computational experiments presented in this paper is to evaluate the performance of the proposed exact Hull Reformulation in comparison to $\varepsilon$-approximation method, which is usually used for any nonlinear constraints. Although the theoretical foundations of the Hull Reformulation have been developed for convex GDPs [16, 11], and most existing literature concentrates on convex GDP formulations![13, 11, 7], as highlighted above, Hull Reformulation is a valid approach to apply to both convex and non-convex problems. Consequently, this study investigates both convex and non-convex GDPs to provide a comprehensive comparison.
@@
The existing literature on non-convex GDPs proposes convexifying the constraints within each disjunct prior to applying techniques such as Hull Reformulation or Big-M to obtain a convex MINLP [21]. However, to the best of our knowledge, no prior work has conducted a comparative analysis between the direct application of the Hull Reformulation and initial constraint convexification. This represents a compelling direction for future research, although it lies beyond the scope of this paper. Similarly, a thorough discussion of the advantages and disadvantages of alternative reformulation methods, such as the Big-M and Binary Multiplication approaches, compared to the Hull Reformulation, especially for non-convex problems, is reserved for future studies. However, the performance results are presented here to provide a comprehensive context.
@@
In this section, we present results of computational experiments for various quadratically constrained GDPs, which were solved using various state-of-the-art general MINLP solvers after applying different reformulation approaches, such as $\varepsilon$-approximation Hull Reformulation, proposed exact Hull Reformulation, Big-M, and Binary Multiplication. All test problems were initially formulated as GDP and reformulated to MINLP using Pyomo (version 6.9.2) [22]. As part of this work, we implemented exact Hull Reformulation for quadratic constraints within Pyomo.GDP [23] by extending its existing Hull Reformulation framework. All computational experiments used the default Pyomo setting of $\varepsilon=10^{-4}$ for $\varepsilon$-approximation approach.
@@
The resulting MINLP formulations were subsequently solved using GAMS (version 49.1.0) [24], employing as sub-solver Gurobi (version 12.0) [25], BARON (version 25.2.1.) [26], and SCIP (version 9.2.0) [27].
All experiments were conducted using a single computational thread on an AMD EPYC 7643 48-Core Processor (96 logical cores), which features 256 MB of L3 cache. The server environment was Red Hat Enterprise Linux 9.5 (Plow), equipped with 1 TB of RAM and 15 GB of swap space.
@@
The following subsections summarize the main findings from the computational experiments across different classes of quadratically constrained GDP problems. These include comparisons between multiple solvers and reformulation strategies. Detailed numerical results and problem generation scripts are available online. [^2] Available at: https://github.com/SECQUOIA/exact_quadratic_hull.
6.1 Random Instances
@@
To evaluate the effectiveness of the proposed Hull Reformulation comprehensively, we conducted computational experiments using randomly generated instances of quadratically constrained GDPs, with quadratic disjunctive constraints, no global constraints, and a quadratic objective function. These problems featured dense matrices and vectors of coefficients, meaning all coefficients were populated with nonzero values. Coefficients were randomly generated from a uniform distribution over the interval [1,1][-1,1]. For the convex problem instances, eigenvalues of the generated random matrices $Q$ were adjusted to be non-negative to ensure positive semidefiniteness of $Q$, thereby guaranteeing convexity of the problem.
@@
Figures 2 and 3 show the cumulative number of instances solved within a given wall-clock time for the convex and non-convex random GDP instances, respectively. As illustrated, the proposed exact Hull Reformulation consistently outperforms the $\varepsilon$-approximation approach, solving substantially more instances within the same time.
It is also important to address a significant issue concerning the number of problems solved successfully, particularly with the $\varepsilon$-approximation method. In the conducted computational experiments, the solvers often erroneously identified solutions as optimal within the specified tolerance $(10^{-6})$ to the lower bound or incorrectly labeled instances as infeasible, despite the explicit feasibility of all generated problems. However, for the same problems, other reformulations or alternative solvers identified feasible solutions with significantly better objective values. We attribute these inconsistencies primarily to numerical instabilities associated with the $\varepsilon$-approximation method.
Table 1 summarizes the number of instances in which the solvers returned incorrect results. A solution was considered incorrect if the solver declared the instance infeasible or, without reaching the time limit, reported an optimal solution with objective value relatively larger by more than $10^{-4}$ compared to those obtained by other reformulations, despite the solver’s optimality tolerance being set to $10^{-6}$. The threshold of $10^{-4}$ was chosen to ensure that such cases represent definitively incorrect solutions rather than results only marginally exceeding the specified tolerance. These incorrect solutions were excluded from the performance profile data shown in Figures 2 and 3. This filtering, for example, explains the absence of a line for the $\varepsilon$-approximation in Figure 3(b), since although it solved 81 instances within the time limit, none of the solutions were correct.
Additionally, in the case of SCIP, the solver sometimes reported an error indicating that the problem failed to converge due to numerical instability, rather than providing a result. This is the reason why, in some cases, the total number of reported solutions in Table 1 is less than the total number of tested instances.
Results presented in Table 1 show that the $\varepsilon$-approximation method had the highest frequency of wrong results in tested problems, and the proposed exact Hull Reformulation demonstrated better numerical stability and reported wrong solutions in fewer cases.
In addition to performance plots, Figures 4 and 5 provide direct comparisons of solution times for the same problems solved using the proposed exact Hull Reformulation and the $\varepsilon$-approximation approach. These figures indicate that whenever both methods successfully converged to the same correct solution, the exact Hull Reformulation achieved faster solutions in the vast majority of cases.
In summary, computational experiments on randomly generated quadratically constrained GDPs demonstrate that the proposed exact Hull Reformulation offers not only improved computational efficiency but also enhanced numerical stability, resulting in fewer solver failures and fewer incorrect solutions compared to the $\varepsilon$-approximation method.
@@
Figure 4: Solution time for random convex quadratically constrained GDPs using Gurobi, BARON, and SCIP for the same problems, comparing the $\varepsilon$-approximation and the proposed exact Hull reformulation.
(a) Gurobi
(b) BARON
(c) SCIP
Figure 5: Solution time for random non-convex quadratically constrained GDPs using Gurobi, BARON, and SCIP for the same problems, comparing the $\varepsilon$-approximation and the proposed exact Hull reformulation.
Table 1: Solver outcomes across reformulation strategies in convex and non-convex cases
(a) Convex Case
Strategy Optimal Timeout Wrong Total
Solver: Gurobi
BigM 240 0 0 240
Hull ($\varepsilon$-approx.) 102 123 15 240
Hull Exact 240 0 0 240
Binary Mult. 211 29 0 240
Solver: BARON
BigM 239 1 0 240
Hull ($\varepsilon$-approx.) 10 127 103 240
Hull Exact 122 88 30 240
Binary Mult. 198 42 0 240
Solver: SCIP
BigM 240 0 0 240
Hull ($\varepsilon$-approx.) 57 168 7 232
Hull Exact 140 95 0 235
Binary Mult. 112 127 0 239
(b) Non-convex Case
Strategy Optimal Timeout Wrong Total
Solver: Gurobi
BigM 97 3 0 100
Hull ($\varepsilon$-approx.) 47 27 26 100
Hull Exact 81 19 0 100
Binary Mult. 91 9 0 100
Solver: BARON
BigM 100 0 0 100
Hull ($\varepsilon$-approx.) 0 19 81 100
Hull Exact 38 62 0 100
Binary Mult. 98 2 0 100
Solver: SCIP
BigM 94 6 0 100
Hull ($\varepsilon$-approx.) 16 80 0 96
Hull Exact 46 45 0 91
Binary Mult. 52 48 0 100
6.2 Continuously Stirred Tank Reactor network problem
Consider the following superstructure optimization problem involving a network of Continuously Stirred Tank Reactors (CSTRs). The objective is to minimize the total volume of CSTRs in series, in which the autocatalytic reaction $A+B\rightarrow 2B$ takes place. The superstructure allows a recycle stream originating from the last reactor to be returned to any preceding reactor within the series. An illustration of this problem is provided in Figure 6. The details and formulations of the problem were previously described in the literature [28, 9] and are summarized in Appendix C.
Originally, the constraints defining this optimization problem were linear or cubic. Since any polynomial constraint can be systematically represented using quadratic constraints, for the purposes of this work, cubic constraints are reformulated into equivalent quadratic forms, using quadratic equality constraints. Consequently, due to the presence of quadratic equality constraints, the resulting GDP formulation is inherently non-convex. Therefore, this problem allows for testing reformulations in an application, where constraints are linear and non-convex quadratic. This structure makes the problem particularly suitable for evaluating reformulation strategies in applications where constraints are both linear and non-convex quadratic.
Figure 6: Superstructure representation for the continuously stirred tank reactor (CSTR) network problem. The configuration includes $N_{T}$ reactor stages arranged in series, where each stage can either be a reactor or a bypass. The flow proceeds from left to right, and a recycle stream from the last reactor is allowed to return to any upstream stage. This flexible structure enables the optimizer to choose the optimal number of reactors and location of recycle to minimize the total reactor volume. [28, 9].
Computational experiments were conducted to evaluate solution times for the CSTR network optimization problem under various reformulation strategies. In these experiments, the maximum permissible number of reactors in the network was varied, keeping all other model inputs unchanged.
Figure 7 summarizes the computational performance obtained in this work. The clearest difference between reformulations was seen when problems were solved using BARON. As shown in Figure 7(b), the exact Hull Reformulation performed similarly or notably better for all problems in terms of solution time compared to the $\varepsilon$-approximation when solved using BARON. In several instances, the proposed exact formulation provided faster solutions compared to all other methods tested.
(a) Gurobi
(b) BARON
(c) SCIP
Figure 7: Solve time comparison of different reformulations of GDP for the CSTR network problem using Gurobi, BARON, and SCIP.
The results obtained using Gurobi showed a similar trend but with less distinct advantages. Although the exact Hull Reformulation still solved the majority of instances more rapidly compared to $\varepsilon$-approximation, there were cases where the $\varepsilon$-approximation slightly outperformed the exact reformulation. However, the differences in solution times between these reformulations (or in fact between any reformulation tested) were generally minor. These results position the exact Hull Reformulation as a competitive and reliable option for solving this problem with Gurobi, performing better or on par with the $\varepsilon$-approximation while offering theoretical advantages.
Figure 7(c) presents the results obtained using SCIP. For a given reformulation, the solution times fluctuated by orders of magnitude as the number of reactors increased. This behavior indicates high variability in solution time in response to small modifications. Increasing the number of reactors results in a systematic and gradual increase in constraints and variables, while preserving the problem structure, and thus, there should be a corresponding gradual increase in solution time. However, this expected trend was not observed in this case. Therefore, these results do not permit a meaningful comparison between reformulation strategies when using SCIP on this problem.
6.3 $k$-Means Clustering
@@
Figure 8 presents the cumulative number of instances solved as a function of wall‑clock time for each reformulation across the three solvers. In many instances, particularly those solved with SCIP, the time limit was reached, so the profiles are supplemented with cumulative plots showing the absolute optimality gap (the difference between the upper and lower bounds) for these cases. Figure 9 provides a direct comparison of solution times for the same instances across reformulations. A summary of solver outcomes is provided in Table 2.
Across all solvers, the proposed exact Hull Reformulation consistently showed improved numerical stability and outperformed the $\varepsilon$-approximation in terms of the number of instances solved correctly within the time limit. For SCIP and Gurobi, most instances were solved faster with the proposed approach. However, for BARON, the $\varepsilon$-approximation achieved faster solve times for the subset of instances where both approaches converged to the correct solution (Figure 9(b)). However, only 38 out of 96 instances solved using the $\varepsilon$-approximation yielded correct solutions within the time limit, compared to 64 for the proposed reformulation.
We hypothesize that the improved performance of the $\varepsilon$-approximation for this particular problem in BARON may be due to internal convexity detection or solver-specific optimizations tailored to the conventional reformulation and this problem type. Nonetheless, even for BARON from an overall standpoint, the proposed reformulation produced fewer incorrect solutions, and thus more problems solved in the given time limit, affirming its robustness and superior practical value.
These results further demonstrate the advantages of the proposed method for convex GDP formulations involving quadratic constraints such as those arising in $k$-means clustering.
@@
(a) Gurobi
(b) BARON
(c) SCIP
Figure 9: Solution time for $k$-means clustering problem using Gurobi, BARON, and SCIP for the same problems, comparing the $\varepsilon$-approximation and the proposed exact Hull reformulation.
@@
Strategy Optimal Timeout Wrong Total
@@
BigM 84 12 0 96
Hull ($\varepsilon$-approx.) 34 62 0 96
Hull Exact 82 14 0 96
Binary Mult. 40 56 0 96
@@
BigM 91 5 0 96
Hull ($\varepsilon$-approx.) 38 1 57 96
Hull Exact 64 32 0 96
Binary Mult. 74 22 0 96
@@
BigM 19 77 0 96
Hull ($\varepsilon$-approx.) 1 95 0 96
Hull Exact 13 83 0 96
Binary Mult. 17 79 0 96
6.4 Constrained layout problem

### v2

@@ 4 Computational experiments
The primary objective of the computational experiments presented in this paper is to evaluate the performance of the proposed GEHR and CEHR (for convex problems) in comparison to $\varepsilon$-approximation method, which is usually used for any nonlinear constraints in HR. Although the theoretical foundations of HR have been developed for convex GDPs [5, 10], and most existing literature concentrates on convex GDP formulations [2, 10, 11], as highlighted above, HR is a valid approach to apply to both convex and non-convex problems. Consequently, this study investigates both convex and non-convex GDPs to provide a comprehensive comparison.
@@
The existing literature on non-convex GDPs proposes convexifying the constraints within each disjunct prior to applying techniques developed for convex GDPs, such as HR or Big-M, to obtain a convex MINLP [25, 30]. A comparison between this approach and the direct application of HR or Big-M reformulation to non-convex MINLPs represents a compelling direction for future research, although it lies beyond the scope of this paper.
Similarly, a detailed assessment of the relative advantages and disadvantages of alternative reformulation strategies, such as Big-M or Binary Multiplication, compared to HR, particularly in the context of nonconvex problems, is deferred to future studies. Performance results for the non-hull reformulations (Big-M and Binary Multiplication) are reported as obtained directly from Pyomo.GDP (i.e., without Big-M parameter tuning) [7]. These results are included solely to provide a broader context. We emphasize that a comprehensive comparison of the proposed hull reformulations with other reformulation methods is not the objective of this work.
@@
In this section, we present results from computational experiments on various quadratically constrained GDPs, solved using state-of-the-art general MINLP solvers after applying different reformulation approaches. All test problems were initially formulated as GDP and reformulated to MINLP using Pyomo (version 6.9.2) [4]. As part of this work, we implemented GEHR and CEHR for quadratic constraints within Pyomo.GDP [7] by extending its existing HR framework. All computational experiments in the main part of the paper used the default Pyomo setting of $\varepsilon=10^{-4}$ for $\varepsilon$-approximation approach, since the values of $\varepsilon$ were found not to influence results qualitatively. For details on the effect of different values of $\varepsilon$, refer to Appendix D.
@@
The resulting MINLP formulations were subsequently solved in GAMS (version 52.2.0) [9] using Gurobi (version 13.0) [15], BARON (version 25.11.17.) [28], and SCIP (version 9.2.4) [3].
All problems were solved with a wall-clock time limit of 5 minutes (300 seconds).
To reduce the influence of solver stopping criteria on the reported outcomes across solvers, we attempted to align key tolerance parameters as closely as possible, including feasibility tolerances, integrality tolerances, and optimality gap tolerances. In our implementation, we used a common target set of tolerances (relative MIP gap $10^{-6}$, absolute MIP gap $10^{-10}$, feasibility tolerance $10^{-6}$, optimality tolerance $10^{-6}$, and integrality tolerance $10^{-5}$) and mapped them to the closest available options in each solver interface.
Nevertheless, exact matching of numerical settings across Gurobi, SCIP, and BARON is not always possible because solvers expose different parameterizations and may interpret tolerances differently. Importantly, this work is not intended to compare solvers against each other; rather, it compares reformulation strategies within each solver. Consequently, the reported data should not be used to draw conclusions about the relative performance of the solvers themselves.
All experiments were run single-threaded on an AMD EPYC 7643 CPU with 1 TB RAM, under Red Hat Enterprise Linux 9.5. Solver parallelism was disabled by setting the solver thread count to 1.
@@
The following subsections summarize the main findings from computational experiments on different quadratically constrained GDP problems. These experiments compare reformulation strategies using different solvers. Detailed numerical results and problem generation scripts are available online.  [^2] Available at: https://github.com/SECQUOIA/exact_quadratic_hull.
4.1 Random Instances (convex and non-convex GDPs)
@@
To comprehensively evaluate the effectiveness of the proposed Hull Reformulations, we conducted computational experiments on randomly generated instances of quadratically constrained GDPs with quadratic disjunctive constraints, no global constraints, and a quadratic objective function. These problems featured dense matrices and vectors of coefficients, meaning all coefficients were nonzero. Coefficients were randomly generated from a uniform distribution over the interval [1,1][-1,1], and all continuous variables were bounded by the interval [1,1][-1,1]. For the convex problem instances, eigenvalues of the generated random matrices $Q$ were adjusted to be non-negative to ensure positive semidefiniteness of $Q$, thereby guaranteeing convexity of the problem.
@@
Figures 2 and 3 report cumulative performance profiles (instances solved versus wall-clock time) for the convex and non-convex problems, respectively. Table 1 summarizes solver outcomes, including timeouts and cases where a solver returned an infeasibility declaration (although the problems were created that ensured that they are feasible) or an “Objective mismatch” solution (these are treated as incorrect and excluded from the profile curves). A solution was classified as “Objective mismatch” if the solver, without reaching the time limit, reported an optimal solution with an objective value relatively larger by more than $10^{-4}$ compared to those obtained by other reformulations, despite the solver’s optimality tolerance being set to $10^{-6}$. The threshold of $10^{-4}$ was chosen to ensure that such cases represent definitively incorrect solutions rather than results only marginally exceeding the specified tolerance.
Objective mismatch arises from an incorrect (overly large) reported lower bound, which can cause the solver to certify a suboptimal incumbent as optimal. To illustrate this, Table 2 reports the objective and solver-reported lower bound on one random convex quadratic instance. As shown in Table 2, solvers sometimes reported matching lower and upper bounds; however, other reformulations or solvers with the same reformulation reported significantly better solutions.
@@
Table 1: Solver outcomes across reformulation strategies in convex and non-convex cases for random quadratic GDPs
(a) Convex case
Strategy Optimal Timeout Infeasible Objective mismatch Total
Solver: Gurobi
BigM 240 0 0 0 240
Hull $\varepsilon$-approx. ($\varepsilon=10^{-4}$) 2 0 200 38 240
General Exact Hull 240 0 0 0 240
CEHR 240 0 0 0 240
Binary Mult. 167 73 0 0 240
Solver: BARON
BigM 233 7 0 0 240
Hull $\varepsilon$-approx. ($\varepsilon=10^{-4}$) 5 233 0 2 240
GEHR 29 183 0 28 240
CEHR 20 190 0 30 240
Binary Mult. 180 60 0 0 240
Solver: SCIP
BigM 240 0 0 0 240
Hull $\varepsilon$-approx. ($\varepsilon=10^{-4}$) 47 182 0 11 240
GEHR 104 136 0 0 240
CEHR 240 0 0 0 240
Binary Mult. 94 146 0 0 240
Solver: SCIP (convex flag)
BigM 240 0 0 0 240
Hull $\varepsilon$-approx. ($\varepsilon=10^{-4}$) 238 0 0 2 240
GEHR 4 0 217 19 240
CEHR 240 0 0 0 240
Binary Mult. 0 0 219 21 240
(b) Non-convex case
Strategy Optimal Timeout Infeasible Objective mismatch Total
Solver: Gurobi
BigM 91 9 0 0 100
Hull $\varepsilon$-approx. ($\varepsilon=10^{-4}$) 2 86 0 12 100
GEHR 59 41 0 0 100
Binary Mult. 83 17 0 0 100
Solver: BARON
BigM 95 4 0 1 100
Hull $\varepsilon$-approx. ($\varepsilon=10^{-4}$) 0 99 1 0 100
GEHR 7 86 0 7 100
Binary Mult. 88 12 0 0 100
Solver: SCIP
BigM 91 9 0 0 100
Hull $\varepsilon$-approx. ($\varepsilon=10^{-4}$) 11 89 0 0 100
GEHR 41 59 0 0 100
Binary Mult. 39 61 0 0 100
Table 2: Objective value, solver-reported lower bound, and wall-clock time for all reformulations on a random convex quadratic instance with parameters $|K|=3,\;|D_{k}|=10,\;\;n=4$. Objective mismatch corresponds to cases where a solver returns optimal, but the reported lower bound is inconsistent with the best objective values obtained by other reformulations on the same instance, indicating an invalid lower bound.
Strategy Time (s) Objective Lower bound
Solver: Gurobi
BigM 0.7864 -1.503530467 -1.503531622
Hull $\varepsilon$-approx. ($\varepsilon=10^{-4}$) 29.6530 -1.209200146 -1.209200146
GEHR 1.2564 -1.503530467 -1.503530784
CEHR 1.3155 -1.503530107 -1.503530472
Binary Mult. 3.3902 -1.503531523 -1.503536828
Solver: BARON
BigM 1.8185 -1.503530473 -1.503530415
Hull $\varepsilon$-approx. ($\varepsilon=10^{-4}$) 301.7905 -1.382069355 -2.172538773
GEHR 196.1422 -1.439272090 -1.439272090
CEHR 279.6740 -1.451448855 -1.451448859
Binary Mult. 4.9766 -1.503530473 -1.503530490
Solver: SCIP
BigM 4.9693 -1.503530165 -1.503530994
Hull $\varepsilon$-approx. ($\varepsilon=10^{-4}$) 4.7891 -1.070302464 -1.070302474
GEHR 27.2644 -1.503530585 -1.503531213
CEHR 8.4097 -1.503530345 -1.503530886
Binary Mult. 23.7996 -1.503530282 -1.503531068
Solver: SCIP (convex flag)
BigM 4.9459 -1.503530165 -1.503530994
Hull $\varepsilon$-approx. ($\varepsilon=10^{-4}$) 4.0870 -1.503530309 -1.503530995
GEHR 0.9600 n/a $\infty$
CEHR 8.3809 -1.503530345 -1.503530886
Binary Mult. 0.8000 n/a $\infty$
Failures to produce a correct solution, whether by incorrectly declaring the problem infeasible or by reporting a solution as optimal with an incorrect lower bound, are attributed to numerical instabilities, which can arguably be particularly pronounced when using the $\varepsilon$-approximation.
4.1.1 Convex instances
As can be seen from the results on the convex benchmark set presented in Figure 2 and Table 1(a), CEHR is, in most cases, the fastest and most reliable HR approach across solvers, with the clearest gains for solvers that recognize conic structure (Gurobi and SCIP). This outcome is consistent with the formulation’s design: it preserves the exact hull relaxation while expressing the perspective term via a rotated-cone-representable constraint, enabling the solver to certify and exploit convexity directly.
In BARON, all hull formulations performed relatively poorly overall. Nevertheless, CEHR and GEHR showed similar performance and were notably better than the $\varepsilon$-approximation in terms of runtime. Both exact formulations exhibited an increased number of “Objective mismatch” outcomes; however, interpreting numerical robustness for the $\varepsilon$-approximation is difficult in BARON because it timed out on 233 of 240 instances in the case of $\varepsilon$-approximation, leaving too few completed runs to support a meaningful comparison.
Finally, GEHR substantially improved solution time and numerical robustness relative to the $\varepsilon$-approximation across all solvers, but it was significantly outperformed by CEHR for solvers with conic recognition (Gurobi and SCIP).
To isolate the role of convexity recognition, we additionally tested the convex random instances with SCIP under a setting that forces SCIP to assume that all constraints are mixed-integer convex functions (reported as “convex flag” results in Table 1(a)).
When SCIP is guided to treat the constraints as mixed-integer convex, the $\varepsilon$-approximation hull behaves much more like CEHR in terms of speed and reliability. This suggests that only with explicit guidance, the branch-and-bound performance of these two formulations is comparable, and they target essentially the same convex relaxation. However, without explicit guidance, the solvers tested here did not reliably recognize or exploit the convexity of the $\varepsilon$-approximation constraints on these instances, which largely explains its substantially worse performance relative to CEHR.
Additionally, the fact that most solutions were incorrect for GEHR under the convex-flag setting underscores its main limitation. Although the feasible region of the relaxation is convex, the reformulated inequality is not necessarily convex as a function of the lifted variables. If the solver is forced to treat these constraints as convex, it may generate invalid cuts and return incorrect solutions. Therefore, declaring GEHR as convex is generally inappropriate, even when the underlying relaxation forms a convex feasible region.
Based on these differences in SCIP, we infer that the performance of the $\varepsilon$-approximation could improve substantially when solvers correctly recognize and exploit the convexity of the resulting constraints. This highlights a key drawback of the $\varepsilon$-approximation: it complicates the functional structure of constraints and can hinder convexity detection, thereby preventing solvers from fully leveraging convex relaxations.
In practice, it is often impractical, and sometimes impossible, to supply annotations for constraints that arise after GDP reformulation, particularly when solvers do not provide the necessary support (especially if only some constraints are convex). For this reason, we restrict the convexity-annotation experiment to the random convex benchmark set. For the remaining benchmarks, we report results under the default solver behavior, which better reflects typical use cases in which users either do not provide convexity flags or cannot.
4.1.2 Non-convex instances
On the non-convex GDPs, CEHR is not applicable, so we focus on comparing GEHR with the $\varepsilon$-approximation. As shown in Figures 3 and Table 1(b), overall, GEHR outperforms the $\varepsilon$-approximation, yielding shorter solution times across all solvers.
In terms of numerical robustness, the improvement is most pronounced for Gurobi: GEHR produced no incorrect solutions, whereas the $\varepsilon$-approximation returned 12 incorrect solutions. For BARON, GEHR exhibited more solver failures, but once again, this comparison is not very informative because the $\varepsilon$-approximation timed out on 99 out of 100 instances, leaving almost no runs completed within the time limit to assess its numerical stability. For SCIP, no reformulation exhibited numerical issues on this benchmark set.
4.2 Continuously Stirred Tank Reactor network problem (non-convex GDP)
Consider the following superstructure optimization problem involving a network of Continuously Stirred Tank Reactors (CSTRs). The objective is to minimize the total volume of CSTRs in series, in which the autocatalytic reaction $A+B\rightarrow 2B$ takes place. The superstructure allows a recycle stream originating from the last reactor to be returned to any preceding reactor within the series. An illustration of this problem is provided in Figure 4. The details and formulations of the problem were previously described in the literature [21] and are summarized in Appendix E.
Originally, the constraints defining this optimization problem were linear or cubic. Since any polynomial constraint can be systematically represented using quadratic constraints, for the purposes of this work, cubic constraints are reformulated into equivalent quadratic forms, using quadratic equality constraints. Consequently, due to the presence of quadratic equality constraints, the resulting GDP formulation is inherently non-convex. Therefore, this problem allows for testing reformulations in an application, where constraints are linear and non-convex quadratic.
Figure 4: Superstructure representation for the continuously stirred tank reactor (CSTR) network problem. The configuration includes $N_{T}$ reactor stages arranged in series, where each stage can either be a reactor or a bypass. The flow proceeds from left to right, and a recycle stream from the last reactor is allowed to return to any upstream stage. This flexible structure enables the optimizer to choose the optimal number of reactors and the location of recycle to minimize the total reactor volume [21].
To vary problem size and difficulty, we changed the maximum allowable number of reactor stages while keeping all other model parameters fixed. Figure 5 summarizes the results across reformulations and solvers. No numerical issues were observed for this benchmark across the tested solvers and reformulations, so we omit a table of solution outcomes.
@@ 4.2 Continuously Stirred Tank Reactor network problem (non-convex GDP)
Figure 5: Cumulative number of instances solved versus solution time for different reformulations applied to the CSTR network problem formulated as GDP. Results are shown separately for Gurobi, BARON, and SCIP. Each curve represents a distinct reformulation strategy.
GEHR performed similarly to the $\varepsilon$-approximation in SCIP and was solved clearly faster in Gurobi and BARON, demonstrating an overall advantage for GEHR.
Although a detailed comparison between Big-M and HRs for non-convex GDPs is outside the scope of this paper, it is worth noting that, for SCIP, both HRs were substantially faster than the Big-M reformulation. For BARON, the $\varepsilon$-approximation performed similarly to Big-M, while GEHR was clearly faster. These results demonstrate that a systematic study of Big-M versus HRs applied to non-convex GDPs could be a valuable direction for future work.
4.3 k-Means Clustering (convex GDP)
@@
Figure 6 presents cumulative performance profiles, extended by the information on the absolute optimality gap. A summary of the solver outcomes is provided in Table 3.
Consistent with the random benchmark results, CEHR delivered the best overall performance on the $k$-means clustering instances across all tested solvers. In addition, no numerical issues were observed for CEHR in any solver.
For Gurobi and SCIP, GEHR performed nearly identically to CEHR and significantly better than $\varepsilon$-approximation. This indicates that, for this problem class, the solvers treat the two formulations similarly. A likely explanation is that, unlike the dense random instances, GEHR constraints arising in $k$-means are recognized as convex or conic-representable without explicit lifting, rather than being handled as generic non-convex quadratic expressions.
For BARON, the $\varepsilon$-approximation achieved slightly faster solve times than GEHR on the subset of instances where both reformulations produced correct solutions, although CEHR remained substantially superior overall. However, GEHR was markedly more numerically robust than the $\varepsilon$-approximation in BARON: the $\varepsilon$-approximation produced 66 incorrect outcomes (infeasible declarations or incorrect optimal solutions), compared to 13 for GEHR.
@@
Strategy Optimal Timeout Infeasible Objective mismatch Total
@@
BigM 76 20 0 0 96
Hull($\varepsilon$-approx.) 26 70 0 0 96
GEHR 69 27 0 0 96
CEHR 66 30 0 0 96
Binary Mult. 28 68 0 0 96
@@
BigM 83 13 0 0 96
Hull($\varepsilon$-approx.) 27 3 18 48 96
GEHR 35 48 0 13 96
CEHR 61 35 0 0 96
Binary Mult. 59 37 0 0 96
@@
BigM 69 27 0 0 96
Hull($\varepsilon$-approx.) 7 88 0 1 96
GEHR 45 51 0 0 96
CEHR 42 54 0 0 96
Binary Mult. 53 43 0 0 96
4.4 Constrained layout problem (convex GDP)